if electric potential at a point in space is given by v=A(xy+yz +zx) and v at (2,-1,8) is 120 volts find electric potential at a point (1,2,3).

Grade:11

2 Answers

Sukhendra Reddy Rompally B.Tech Mining Machinery Engg, ISM Dhanbad
13 years ago

just substitute x,y,z in the eqn and u get A=20

now dat A=20,again substitute x,y,z = 1,2,3 resp and u get v=20*11= 220

plzzz approve my answer if u like it

AJIT AskiitiansExpert-IITD
13 years ago

DEAR NIKITHA ,

 

simply put the values

120  =A( 2*(-1) + (-1) *8 + 8 *(2))

120  = A(6)

=>  A = 20.

 

so again for (1,2,3) 

v  =  20 (1*2 + 2*3 + 3* 1)

v = 220v

GUD LUCK !!!!


Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian